# E2-BKI: Evidential Ellipsoidal Bayesian Kernel Inference for Uncertainty-aware Gaussian Semantic Mapping

### v1

@@ Abstract
Semantic mapping aims to construct a 3D semantic representation of the environment, providing essential knowledge for robots operating in complex outdoor settings. While Bayesian Kernel Inference (BKI) addresses discontinuities of map inference from sparse sensor data, existing semantic mapping methods suffer from various sources of uncertainties in challenging outdoor environments. To address these issues, we propose an uncertainty-aware semantic mapping framework that handles multiple sources of uncertainties, which significantly degrade mapping performance. Our method estimates uncertainties in semantic predictions using Evidential Deep Learning and incorporates them into BKI for robust semantic inference. It further aggregates noisy observations into coherent Gaussian representations to mitigate the impact of unreliable points, while employing geometry-aligned kernels that adapt to complex scene structures. These Gaussian primitives effectively fuse local geometric and semantic information, enabling robust, uncertainty-aware mapping in complex outdoor scenarios. Comprehensive evaluation across diverse off-road and urban outdoor environments demonstrates consistent improvements in mapping quality, uncertainty calibration, representational flexibility, and robustness, while maintaining real-time efficiency.
@@ I INTRODUCTION
To address these limitations, we propose an uncertainty-aware semantic mapping framework based on anisotropic Gaussian primitives. Our key insight is that aggregating noisy observations into ellipsoidal primitives enables joint modeling of local geometry, semantics, and uncertainty, while improving robustness to noise. Building on this idea, we construct Gaussian primitives via uncertainty-aware spatial aggregation of neighboring observations, as illustrated in Fig. 1. These primitives undergo refinement through merging and pruning, and form the basis of our evidential ellipsoidal BKI formulation with anisotropic kernels that adapt to geometry and incorporate semantic uncertainty. Extensive experiments across both off-road and urban outdoor environments demonstrate superior performance, consistently outperforming existing methods in semantic accuracy and geometric completeness while maintaining real-time efficiency. In summary, our contributions include:
@@ III Semantic Bayesian Kernel Inference
Figure 2: Key limitations of S-BKI and our solutions. The static isotropic kernel in S-BKI faces three distinct types of uncertainty: (a) semantic uncertainty arising from neural network predictions, (b) spatial uncertainty caused by misalignment with local scene geometry, and (c) observation uncertainty from processing each observation independently, compromising consistency. Our method addresses these by (a) prioritizing reliable information via uncertainty estimates, (b) adapting kernels to local scene geometry, and (c) leveraging local context through Gaussian primitives.
@@ III-B Limitations of S-BKI Framework
While S-BKI enables continuous semantic mapping, it suffers from three fundamental limitations in uncertainty handling that compromise reliability in complex environments. We highlight these gaps and propose corresponding solutions, as illustrated in Fig. 2.
@@ III-B2 Spatial Uncertainty
The isotropic kernel in (2) introduces spatial uncertainty by applying uniform support across all directions, regardless of local scene geometry. This uniform support fails to account for the anisotropic nature of real-world structures. When processing elongated structures like roads or fences, spherical kernels propagate semantic information orthogonally to their structural directions, causing geometric misalignment and blurring semantic boundaries, as illustrated in Fig. 1. Moreover, the fixed radius $\ell$ cannot adapt to varying point densities across the scene, leading to excessive smoothing in dense regions and insufficient coverage in sparse areas. Our framework addresses these issues through geometry-aligned anisotropic kernels that adapt their spatial support to local scene structure, steering influence along principal geometric directions while providing adaptive coverage based on local point density (Fig. 2).
@@ III-B3 Observation Uncertainty
We define observation uncertainty as the inherent unreliability arising from sensor noise, limited resolution, sparse point coverage, and calibration errors. Although the update rule (3) incorporates neighboring points for the posterior update, each point is processed independently without leveraging local context. This isolation makes the system susceptible to noise, as individual errors directly affect the semantic map without being corrected by nearby consistent observations. As illustrated in Fig. 3, this susceptibility is further exacerbated with increasing sensor distance, where points become sparser and less informative while calibration errors compound the unreliability of distant observations. These limitations motivate our design of spatially coherent primitives that capture local context, enabling noise reduction via local consistency while providing a more robust foundation for semantic inference (Fig. 2).
Figure 3: Observation uncertainty varies with sensor distance. Proximal regions benefit from dense, high-resolution data that enables reliable inference, while distal regions suffer from sparse coverage and reduced visual detail, leading to higher observation uncertainty.
@@ IV-A Evidential Semantic Segmentation
To quantify uncertainties of semantic predictions, we adopt Evidential Deep Learning (EDL) [22] to extend one-hot semantic predictions $\mathbf{y}_{n}$ into semantic probability $\mathbf{p}_{n}$ with uncertainty estimates $u_{n}$, extending the input $\mathcal{D}$ to evidential points $\mathcal{D^{\prime}}=\{(\mathbf{x}_{n},\mathbf{p}_{n},u_{n})\}^{N}_{n=1}$. This formulation follows EBS [10], which presents EDL-based uncertainty estimates of semantic prediction and its integration into uncertainty-aware semantic mapping. Based on Dempster-Shafer Theory of evidence (DST) [23], EDL estimates per-class evidence in a single forward pass and converts it into class probability $\mathbf{p}_{n}$ and corresponding uncertainty $u_{n}$. These uncertainty estimates reflect the model’s confidence in its predictions, which helps reduce the impact of uncertain predictions during fusion (see [10] for details). The evidential points are then aggregated into Gaussian primitives through uncertainty-aware fusion to construct robust semantic representations.
@@ IV-B Gaussian Initialization
To this end, we begin by clustering the evidential points into $J$ clusters $\{\mathcal{C}_{j}\}_{j=1}^{J}$ using the K-Means++ algorithm [24]. Each cluster $\mathcal{C}_{j}$ is then abstracted as a Gaussian primitive $\mathcal{G}_{j}=(\mathbf{g}_{j},\mathbf{p}_{j},u_{j})$, where the geometric component $\mathbf{g}_{j}=(\bm{\mu}_{j},\bm{\Sigma}_{j})$ models the local geometry through mean $\bm{\mu}_{j}$ and covariance $\bm{\Sigma}_{j}$. The semantic probability $\mathbf{p}_{j}$ and uncertainty $u_{j}$ are derived by aggregating the semantic predictions $(\mathbf{p}_{n},u_{n})$ from all points in $\mathcal{C}_{j}$ through DST-based combination rules [25]. Specifically, we convert each probability $\mathbf{p}_{n}$ into belief mass $\mathbf{b}_{n}$ with $b_{n}^{c}=p_{n}^{c}-u_{n}/C$, and apply the combination rule for combining two belief masses:
$b^{c}=\frac{1}{1-\delta}(b_{1}^{c}b_{2}^{c}+b_{1}^{c}u_{2}+b_{2}^{c}u_{1}),\ \ u=\frac{1}{1-\delta}u_{1}u_{2},$ (5)
where $C$ is the number of semantic classes, and $\delta=\sum_{x\neq y}b_{1}^{x}b_{2}^{y}$ is a measure of the conflict between two belief masses. We iteratively apply this pairwise combination rule to fuse all $|\mathcal{C}_{j}|$ belief masses within the cluster, progressively accumulating evidence. Then, the merged belief mass $\mathbf{b}_{j}$ is converted back into the probability via $p_{j}^{c}=b_{j}^{c}+u_{j}/C$. This uncertainty-aware fusion effectively resolves semantic conflicts within each cluster and yields robust semantics with reliable uncertainty estimates (for more details, see [11]).
@@ IV-C1 Merging
As BKI incrementally accumulates evidence over time (Section III-A), Gaussian primitives similarly accumulate through iterative initialization. This temporal accumulation can cause redundancy from overlapping regions, while some primitives may consist of insufficient points for stable geometric estimation. To address these issues, we merge nearby primitives with the same semantic label. Geometric components $\mathbf{g}_{j}$ are merged by combining their statistical moments, while semantic components $(\mathbf{p}_{j},u_{j})$ are fused through DST-based combination rules as in (5). This merging stabilizes both geometric and semantic representations and reduces computational overhead by decreasing the total number of primitives.
@@ IV-D Evidential Ellipsoidal Kernel
For geometry-aligned kernel support, we compute the distance $\mathbf{d}(\mathbf{\hat{x}}_{m},\mathcal{G}_{j})$ by representing each primitive as an anisotropic ellipsoid. We then find the closest point on the ellipsoid surface to query point $\mathbf{\hat{x}}_{m}$ through the following minimization:
$\min_{\mathbf{v}\in\mathbb{R}^{3}}\|\mathbf{\hat{x}}_{m}-\mathbf{v}\|^{2}\ \text{s.t.}\ (\mathbf{v}-\bm{\mu}_{j})^{\top}{\bm{\Sigma}_{j}}^{-1}(\mathbf{v}-\bm{\mu}_{j})=\tau,$ (7)
where $\tau$ controls the ellipsoid size and the resulting minimum $d^{*}:=\mathbf{d}(\mathbf{\hat{x}}_{m},\mathcal{G}_{j})$ used in our kernel. For queries inside the ellipsoid, we set $\mathbf{d}=0$. This formulation enables kernels to conform to elongated structures and propagate semantic information along principal geometric directions, addressing the spatial uncertainty of isotropic kernels.
@@
where $\beta$ controls uncertainty sensitivity and $\mathrm{U}_{\text{thr}}$ is dynamically set to exclude the most uncertain $\tilde{u}$ percentile of primitives. This formulation ensures that highly uncertain observations are filtered out while smoothly modulating spatial influence based on semantic uncertainty, enabling robust semantic inference (for more details, see [10]).
@@ V-B1 Implementation Details
For 3D mapping frameworks, we set the voxel resolution to $0.2\textrm{m}$ with hyperparameters in Table I. Gaussian initialization first partitions points by semantic class, then applies K-means++ clustering within each partition. We use 256 clusters for RELLIS-3D and 1024 for other datasets to account for variations in point density. During refinement, we first check for semantic consistency within $d_{L}$ neighborhoods: if all neighboring primitives are semantically consistent, those within $d_{S}$ are merged. When semantic conflicts exist within $d_{L}$ and condition (6) is satisfied, the more distant primitive is pruned. The kernel threshold $\tau$ is chosen to enclose $10\%$ of the probability mass of each Gaussian.
@@ V-B2 Evaluation Metrics
For evaluation, we query the maps at ground truth point coordinates. We adopt metrics established in EBS [10], including $\mathrm{IoU}$ and $\mathrm{mIoU}$ for semantic accuracy with extending the Accuracy to $\mathrm{Acc}=\sum_{c=1}^{C}\text{TP}_{c}/\mathbf{Q}$ for joint geometric and semantic evaluation, where $\text{TP}_{c}$ denotes true positives for class $c$ and $\mathbf{Q}$ is the total number of queries. For uncertainty calibration, we employ Brier Score ($\mathrm{BS}$ $\downarrow$) with normalized $\mathrm{Var}[\hat{\theta}_{m}^{\psi}]$.
@@ V-C2 Uncertainty Calibration
Our framework demonstrates superior uncertainty calibration across all datasets, achieving the best $\mathrm{BS}$ on RELLIS-3D and KITTI-360 while maintaining competitive performance on OffRoad. Fig. 7 shows how our uncertainty-aware processing addresses unreliable semantic predictions with improved uncertainty calibration. While S-BKI produces incorrect semantics with vague uncertainty estimates in ambiguous regions, our method assigns correct labels while assigning high uncertainty values that appropriately flag challenging areas requiring caution. These improvements demonstrate our framework’s ability to deliver accurate semantic maps with reliable uncertainty quantification across diverse environments.
@@ V-D3 Scene Representation Versatility
Figure 7: Semantic and uncertainty map (the third row of Fig. 5). While S-BKI produces incorrect semantics with uncalibrated uncertainty estimates, E2-BKI provides correct semantic labels while appropriately flagging these challenging areas with high uncertainty values, showing improved uncertainty calibration.
@@ V-F Runtime Analysis
Finally, we evaluate the computational efficiency of our framework, averaging over five runs on a laptop with an Intel i7-12700H CPU. Gaussian construction achieves $9.73\pm 0.21$ Hz, and evidential ellipsoidal BKI operates at $12.10\pm 0.23$ Hz. The full framework runs at $5.39\pm 0.11$ Hz, significantly faster than S-BKI baseline ($3.15\pm 0.04$ Hz). This advantage arises from two key factors: (1) abstraction of individual points into compact Gaussians, which reduces the number of BKI computations, and (2) local merging, which minimizes redundant computation, as demonstrated by the throughput drop to $2.70\pm 0.08$ Hz without refinement steps.

### v2

@@ Abstract
Semantic mapping aims to construct a 3D semantic representation of the environment, providing essential knowledge for robots operating in complex outdoor settings. While Bayesian Kernel Inference (BKI) addresses discontinuities of map inference from sparse sensor data, existing semantic mapping methods suffer from various sources of uncertainties in challenging outdoor environments. To address these issues, we propose an uncertainty-aware semantic mapping framework that handles multiple sources of uncertainties, which significantly degrade mapping performance. Our method estimates uncertainties in semantic predictions using Evidential Deep Learning and incorporates them into BKI for robust semantic inference. It further aggregates noisy observations into coherent Gaussian representations to mitigate the impact of unreliable points, while employing geometry-aligned kernels that adapt to complex scene structures. These Gaussian primitives effectively fuse local geometric and semantic information, enabling robust, uncertainty-aware mapping in complex outdoor scenarios. Comprehensive evaluation across diverse off-road and urban outdoor environments demonstrates consistent improvements in mapping quality, uncertainty calibration, representational flexibility, and robustness, while maintaining real-time efficiency. Our project website: https://e2-bki.github.io/
@@ I INTRODUCTION
To address these limitations, we extend the uncertainty-aware semantic mapping framework [17] by introducing anisotropic Gaussian primitives as a compact scene representation. Our key insight is that aggregating noisy observations into ellipsoidal primitives enables joint modeling of local geometry, semantics, and uncertainty, while improving robustness to noise. Building on this idea, we construct Gaussian primitives via uncertainty-aware spatial aggregation of neighboring observations, as illustrated in Fig. 1. These primitives undergo refinement through merging and pruning, and form the basis of our evidential ellipsoidal BKI formulation with anisotropic kernels that adapt to geometry and incorporate semantic uncertainty. Extensive experiments across both off-road and urban outdoor environments demonstrate superior performance, consistently outperforming existing methods in semantic accuracy and geometric completeness while maintaining real-time efficiency. In summary, our contributions include:
@@ III Semantic Bayesian Kernel Inference
Figure 2: Key limitations of S-BKI and our solutions. The static isotropic kernel in S-BKI does not account for (a) semantic uncertainty arising from neural network predictions, (b) spatial uncertainty caused by misalignment with local scene geometry, and (c) observation uncertainty from distant noisy measurements. Our method addresses these by (a) prioritizing reliable information via uncertainty estimates, (b) adapting kernels to local scene geometry, and (c) leveraging local context through Gaussian primitives.
@@ III-B Limitations of S-BKI Framework
While S-BKI enables continuous semantic mapping, its failures in complex environments can be analyzed in terms of three types of uncertainty. We highlight these three aspects and propose corresponding solutions, as illustrated in Fig. 2.
@@ III-B2 Spatial Uncertainty
The isotropic kernel in (2) applies uniform support across all directions, regardless of local scene geometry. We use the term spatial uncertainty to describe the resulting mismatch between this uniform support and the anisotropic nature of real-world structures. When processing elongated structures like roads or fences, spherical kernels propagate semantic information orthogonally to their structural directions, causing geometric misalignment and blurring semantic boundaries, as illustrated in Fig. 1. Moreover, the fixed radius $\ell$ cannot adapt to varying point densities across the scene, leading to excessive smoothing in dense regions and insufficient coverage in sparse areas. Our framework addresses these issues through geometry-aligned anisotropic kernels that adapt their spatial support to local scene structure, steering influence along principal geometric directions while providing adaptive coverage based on local point density (Fig. 2).
@@ III-B3 Observation Uncertainty
We use the term observation uncertainty to describe the inherent unreliability arising from sensor noise, limited resolution, sparse point coverage, and calibration errors. Although the update rule (3) incorporates neighboring points for the posterior update, each point is processed independently without leveraging local context. This isolation makes the system susceptible to noise, as individual errors directly affect the semantic map without being corrected by nearby consistent observations. As illustrated in Fig. 3, this susceptibility is further exacerbated with increasing sensor distance, where points become sparser and less informative while calibration errors compound the unreliability of distant observations. These limitations motivate our design of spatially coherent primitives that capture local context, enabling noise reduction via local consistency while providing a more robust foundation for semantic inference (Fig. 2).
Figure 3: Observation uncertainty varies with sensor distance. Left: sensor-agnostic schematic showing dense, high-resolution proximal observations and sparse, low-detail distal observations. Right: RGB image with semantic labels and LiDAR measurements overlaid; sparser returns at longer ranges lead to higher observation uncertainty.
@@ IV-A Evidential Semantic Segmentation
To quantify uncertainties of semantic predictions, we adopt Evidential Deep Learning (EDL) [28] to extend one-hot semantic predictions $\mathbf{y}_{n}$ into semantic probability $\mathbf{p}_{n}=[p_{n}^{1},\ldots,p_{n}^{C}]$ with uncertainty estimates $u_{n}$, extending the input $\mathcal{D}$ to evidential points $\mathcal{D^{\prime}}=\{(\mathbf{x}_{n},\mathbf{p}_{n},u_{n})\}^{N}_{n=1}$. This formulation follows EBS [17], which presents EDL-based uncertainty estimates of semantic prediction and its integration into uncertainty-aware semantic mapping. Based on Dempster-Shafer Theory of evidence (DST) [30], EDL estimates per-class evidence in a single forward pass and converts it into class probability $\mathbf{p}_{n}$ and corresponding uncertainty $u_{n}$. These uncertainty estimates reflect the model’s confidence in its predictions, which helps reduce the impact of uncertain predictions during fusion (see [17] for details). The evidential points are then aggregated into Gaussian primitives through uncertainty-aware fusion to construct robust semantic representations.
@@ IV-B Gaussian Initialization
To this end, we begin by clustering the evidential points into $J$ clusters $\{\mathcal{C}_{j}\}_{j=1}^{J}$. We use the K-Means++ algorithm [1] as a stable default choice, although other clustering strategies are also compatible with our framework. Each cluster $\mathcal{C}_{j}$ is then abstracted as a Gaussian primitive $\mathcal{G}_{j}=(\bm{\mu}_{j},\bm{\Sigma}_{j},\mathbf{p}_{j},u_{j})$, where the local geometry is modeled by mean $\bm{\mu}_{j}$ and covariance $\bm{\Sigma}_{j}$. The semantic probability $\mathbf{p}_{j}$ and uncertainty $u_{j}$ are derived by aggregating the semantic predictions $(\mathbf{p}_{n},u_{n})$ from all points in $\mathcal{C}_{j}$ through DST-based combination rules [11]. Specifically, we convert each probability $\mathbf{p}_{n}$ into belief mass $\mathbf{b}_{n}$ with $b_{n}^{c}=p_{n}^{c}-u_{n}/C$, and apply the combination rule for combining two belief masses:
$b^{c}=\frac{1}{1-\eta}(b_{1}^{c}b_{2}^{c}+b_{1}^{c}u_{2}+b_{2}^{c}u_{1}),\ \ u=\frac{1}{1-\eta}u_{1}u_{2},$ (5)
where $C$ is the number of semantic classes, and $\eta=\sum_{x\neq y}b_{1}^{x}b_{2}^{y}$ is a measure of the conflict between two belief masses. We iteratively apply this pairwise combination rule to fuse all $|\mathcal{C}_{j}|$ belief masses within the cluster, progressively accumulating evidence. Then, the merged belief mass $\mathbf{b}_{j}$ is converted back into the probability via $p_{j}^{c}=b_{j}^{c}+u_{j}/C$. This uncertainty-aware fusion effectively resolves semantic conflicts within each cluster and yields robust semantics with reliable uncertainty estimates (for more details, see [18]).
@@ IV-C1 Merging
As BKI incrementally accumulates evidence over time (Section III-A), Gaussian primitives similarly accumulate through iterative initialization. This temporal accumulation can cause redundancy from overlapping regions, while some primitives may consist of insufficient points for stable geometric estimation. To address these issues, we merge nearby primitives with the same semantic label. Geometric components $(\bm{\mu}_{j},\bm{\Sigma}_{j})$ are merged by combining their statistical moments, while semantic components $(\mathbf{p}_{j},u_{j})$ are fused through DST-based combination rules as in (5). This merging stabilizes both geometric and semantic representations and reduces computational overhead by decreasing the total number of primitives.
@@ IV-D Evidential Ellipsoidal Kernel
For geometry-aligned kernel support, we compute the distance $d^{\ast}(\mathbf{\hat{x}}_{m},\mathcal{G}_{j})$ by representing each primitive as an anisotropic ellipsoid and finding the closest point $\mathbf{v}^{\ast}$ on the ellipsoid surface to query point $\mathbf{\hat{x}}_{m}$ through the following minimization over surface points $\mathbf{v}$:
$d^{\ast}=\min_{\mathbf{v}\in\mathbb{R}^{3}}\|\mathbf{\hat{x}}_{m}-\mathbf{v}\|^{2}\ \text{s.t.}\ (\mathbf{v}-\bm{\mu}_{j})^{\top}{\bm{\Sigma}_{j}}^{-1}(\mathbf{v}-\bm{\mu}_{j})=\tau,$ (7)
where $\tau$ controls the ellipsoid size and the resulting minimum $d^{\ast}(\mathbf{\hat{x}}_{m},\mathcal{G}_{j}):=d^{\ast}$ used in our kernel. For queries inside the ellipsoid, we set $d^{\ast}(\mathbf{\hat{x}}_{m},\mathcal{G}_{j}):=0$. This formulation enables kernels to conform to elongated structures and propagate semantic information along principal geometric directions, addressing the spatial uncertainty of isotropic kernels.
@@
where $\beta$ controls uncertainty sensitivity and $\mathrm{U}_{\text{thr}}$ is dynamically set to exclude the most uncertain $\tilde{u}$ percentile of primitives. This formulation ensures that highly uncertain observations are filtered out while smoothly modulating spatial influence based on semantic uncertainty, enabling robust semantic inference (for more details, see [17]). When multiple primitives overlap, their kernel-weighted evidence is accumulated in $\bm{\alpha}_{m}$, so primitives with lower uncertainty naturally dominate the posterior semantics while uncertain primitives contribute only weak evidence.
@@ V-B1 Implementation Details
For 3D mapping, we perform a grid search to select the hyperparameters. The chosen values are listed in Table I, and we note that E2-BKI is not sensitive to these values. Gaussian initialization first partitions points by semantic class, then applies K-means++ clustering within each partition. We use 256 clusters for RELLIS-3D and 1024 for other datasets to account for variations in point density. During refinement, we first check for semantic consistency within $d_{L}$ neighborhoods: if all neighboring primitives are semantically consistent, those within $d_{S}$ are merged. When semantic conflicts exist within $d_{L}$ and condition (6) is satisfied, the more distant primitive is pruned. The threshold $\tau$ is chosen to enclose $10\%$ of the probability mass of each Gaussian.
@@ V-B2 Evaluation Metrics
For fair comparison, all methods are evaluated on a $0.2\textrm{ m}$ voxelized grid. We adopt metrics established in EBS [17], including per-class $\mathrm{IoU}$ and $\mathrm{mIoU}$ for semantic accuracy with extending the Accuracy to $\mathrm{Acc}=\sum_{c=1}^{C}\text{TP}_{c}/\mathbf{Q}$ for joint geometric and semantic evaluation, where $\text{TP}_{c}$ denotes true positives for class $c$ and $\mathbf{Q}$ is the total number of queries. For uncertainty calibration, we employ Brier Score ($\mathrm{BS}$ $\downarrow$) with normalized $\mathrm{Var}[\hat{\theta}_{m}^{\psi}]$.
@@ V-C2 Uncertainty Calibration
E2-BKI achieves well-calibrated uncertainty estimates across all datasets, achieving the best $\mathrm{BS}$ on RELLIS-3D and KITTI-360 and competitive performance on OffRoad. Fig. 7 illustrates how our uncertainty-aware processing addresses unreliable semantic predictions. In the boxed region, S-BKI produces incorrect semantics with weakly informative uncertainty, whereas our framework generates more reliable semantic assignments and preserves higher uncertainty in areas where the evidence is insufficient. Consequently, the predicted uncertainty better reflects the difficulty of each prediction and clearly identifies high-risk regions for downstream modules, enabling accurate semantic maps together with well-calibrated and trustworthy uncertainty information across diverse environments.
@@ V-D3 Scene Representation Versatility
Figure 7: Semantic and uncertainty map (the third row of Fig. 5). While S-BKI produces incorrect semantics with uncalibrated uncertainty estimates, E2-BKI recovers the correct semantic labels and explicitly flags these challenging regions with higher uncertainty.
@@ V-F Runtime Analysis
Finally, we evaluate the computational efficiency of our framework, averaging over five runs on a laptop with an Intel i7-12700H CPU. Gaussian construction achieves $9.73\pm 0.21$ Hz, and evidential ellipsoidal BKI operates at $12.10\pm 0.23$ Hz. The full framework runs at $5.39\pm 0.11$ Hz, faster than the S-BKI baseline ($3.15\pm 0.04$ Hz). This advantage arises from (1) the abstraction of individual points into Gaussians, which reduces the number of BKI computations, and (2) local merging, which further reduces redundant computation, as demonstrated by the throughput drop to $2.70\pm 0.08$ Hz without refinement steps.